You sold 52 items and have an income of 2475. How many did you sell for each if the fries are sold for 40, spaghetti for 35, chicken for 60 and you sold 7 more chicken than spaghetti?
Answer by richwmiller(17219)

f+s+c=52,
c=s+7,
40f+35s+60c=2475
c = 24, f = 11, s = 17
check
40*11+35*17+60*24=2475
ok